OpenStudy (anonymous):

simplify the complex fraction.. n-6 ____ n^2+11n+24 _____________ n+1 ____ n+3

OpenStudy (anonymous):

@NNL1991

OpenStudy (anonymous):

@Dido525

OpenStudy (espex):

Recall that, when you divide fractions, you can rewrite the problem as multiplication by the reciprocal. So \(\frac{\frac{a}{b}}{\frac{c}{d}}\) becomes \(\frac{a}{b}*\frac{d}{c}\)

OpenStudy (espex):

You can then factor and you will find that there will be terms to cancel.

OpenStudy (phi):

first step is multiply the bottom fraction by (n+3)/(n+1) (the inverse or "flipped") this makes the bottom 1 you must also multiply the top fraction by (n+3)/(n+1) see eSpex post

OpenStudy (phi):

next, you should factor n^2+11n+24 can you do that ?

OpenStudy (anonymous):

n-6 n+1 what i got so far. ------------ * ---- n^2+11n+24 n+3

OpenStudy (phi):

almost, but it should be (n+3)/(n+1)

OpenStudy (anonymous):

factoring i got( n+8 )(n+3)

OpenStudy (phi):

\[\frac{ \frac{n-6}{n^2+11n+24} \cdot \frac{n+3}{n+1} } { \frac{\cancel{n+1}}{n+3} \cdot \frac{n+3}{\cancel{n+1}}} \]

OpenStudy (phi):

the n+3 in the bottom also cancel, so the complicated bottom becomes 1 and you are left with the top you now have \[ \frac{n-6}{(n+8)(n+3) } \cdot \frac{n+3}{n+1} \] it looks like you can cancel (n+3) from top and bottom

OpenStudy (anonymous):

n-6 ---- n \[\neq-8,-3 or -1\] n^2+9n+8

OpenStudy (phi):

yes, looks good

OpenStudy (anonymous):

whewww :) thanks!

OpenStudy (phi):

the only part that confused you was "flipping" the bottom fraction -- that is the "trick" you should learn . if you see a/b / c/d change it to a/b * d/c

OpenStudy (anonymous):

alright! will do(: thank you!

OpenStudy (anonymous):

phi has done a good job with the answer, so nothing left for me to do. :)